OpenStudy (anonymous):

can (3a-4)(2a²-1) be factored down further ?

OpenStudy (across):

Try and factor \(2a^2-1\) further.

OpenStudy (across):

Use the fact that \(a^2-b^2=(a+b)(a-b)\).

OpenStudy (anonymous):

yes but there is a coefficient

OpenStudy (across):

Factor it out.

OpenStudy (across):

\[2a^2-1=2(a^2-\frac12)\]

OpenStudy (anonymous):

2(3a-4)(a+1)(a-1)

OpenStudy (anonymous):

right ?

OpenStudy (anonymous):

@across

OpenStudy (anonymous):

You can factor it, but not with integers and rational numbers. if you want to break it down into all linear binomials, you can use fractions and irrational numbers.

OpenStudy (anonymous):

Conventionally, no.

Directrix (directrix):

@burhan101 Not over the set of rational numbers. See attached file.